OpenStudy (anonymous):

x + 100 = 651 What is x? Medal and fan

OpenStudy (camper4834):

algebra?

OpenStudy (anonymous):

yep

OpenStudy (nicholausblackmon):

551

OpenStudy (camper4834):

you can subtract 100 from both sides

OpenStudy (anonymous):

Thanks

sammixboo (sammixboo):

Yes, just subtract 100 on both sides of the equation, since you are originally adding x + 100 - 100 = 651 - 100 x = 651 - 100

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

subtract 100 from both sides and you get 551 as the answer

OpenStudy (anonymous):

Thanks Matthew1213 :)

OpenStudy (nicholausblackmon):

WHAT

OpenStudy (nicholausblackmon):

I GAVE ANSWER THO :(

OpenStudy (anonymous):

Thanks NicholausBlackmon for your help :)

OpenStudy (nicholausblackmon):

:)

OpenStudy (anonymous):

Sorry :)

OpenStudy (anonymous):

I wasn't paying attention

OpenStudy (anonymous):

what do medals even do

OpenStudy (anonymous):

Um...

OpenStudy (anonymous):

I dunno, to tell the truth

OpenStudy (nicholausblackmon):

YOU USE THEM TO BUY TITLES

OpenStudy (anonymous):

ok thanks